# E Meena Nanivadekar A Azad Shenoy Roy Chowdhuri Banga Chhabra Bhat Kalluri Chetupalli Vijayasenan Ganapathy

### 03 Mar 2026 v1

@@ Abstract
The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge introduces a conversational AI benchmark focused on understanding goal-oriented, real-world medical dialogues collected in the field. The challenge addresses multi-speaker interactions between healthcare workers and seekers characterized by spontaneous, noisy and overlapping speech across Indian languages and dialects. As part of the challenge, medical conversational dataset comprising 25 hours of development data and 10 hours of blind evaluation recordings was released. We provided baseline systems within a unified end-to-end pipeline across 4 tasks - speaker diarization, automatic speech recognition, topic identification and dialogue summarization - to enable consistent benchmarking. System performance is evaluated using established metrics such as diarization error rate (DER), time-constrained minimum-permutation word error rate (tcpWER), and ROUGE-L. During this evaluation (Phase-I), $12$ teams, across the globe, actively participated pushing the baseline systems on these metrics. However, even with a 686-8 week dedicated effort from various participants, the task is shown to be substantially challenging, and the existing systems are significantly short of healthcare deployment readiness.
@@ 1 Introduction
Speech processing tools for the understanding of conversations from healthcare domain can have transformative effects on public health systems, especially in community healthcare settings. Existing speech datasets in this domain are mostly recorded in controlled environments such as hospitals, involve structured clinician-patient conversations, and are predominantly in English. AI-driven tools developed using such datasets often fail to perform when deployed in real-world setups involving community healthcare workers. Further, there is a lack of research in low-resource languages, in particular in Indian languages. To catalyze the research efforts in this space, we collected and released an annotated corpus of conversations between community health workers and healthcare seekers in informal settings in Hindi. Using the data, we hosted an open leaderboard-style challenge - ``The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge''.
@@ 2 Related Work
In the Indian context, there has been a growing interest in developing ASR systems for low-resource and multilingual settings. However, existing datasets are restricted to isolated utterances or short conversational clips [Bhogale2022EffectivenessOM, bhogale23_interspeech, Nithya2023SPRINGINXAM, javed-etal-2024-indicvoices], which limits their suitability for studying long-form, dialogue centric ASR systems. Further, there are no public datasets covering the conversations in the healthcare domain.
@@
The Phase-I of the DISPLACE-M evaluations was organized between January 1, 2026 to February 27, 2026. The details of the dataset (Section 3), tracks and metrics (Section 4), baseline systems (Section 5) and challenge results (Section 6) are described below.
@@ 3.1 Data Collection
We recorded around $50$ hours of annotated healthcare conversational speech from rural and semi-urban areas, collected through field deployments in the Indian states of Haryana and Bihar. Data collection was conducted by $80$ frontline community health workers across $20$ administrative blocks spanning $10$ districts. The primary participants were NPHWs (non-physician health workers) - ASHA [^1] ASHA (Accredited Social Health Activists) workers, who are government-appointed female community health workers serving as the first point of contact for healthcare seekers (HS) and Anganwadi Sevikas  [^2] Anganwadi sevikas are frontline workers under India’s Integrated Child Development Services (ICDS) program.. Recordings captured natural NPHW–participant conversations occurring during routine community healthcare interactions. These interactions took place in real care environments such as Primary Health Centres (PHCs), Anganwadi centers, primary schools, open village meeting spaces (chaupals), household visits and community outreach settings. In our recordings, Hindi served as the primary language with code-switching to Indian English and with heavy use of regional dialects, including Haryanvi, Bhojpuri, and Magahi, reflecting natural linguistic practices.
The recorded conversations include speakers from different age groups. Each recording ranges from 55-$30$ minutes, typically involving goal-oriented conversations between two participants - NPHW and a healthcare seeker. The conversations, which include nature field ambiance, primarily focused on topics such as General Health and Constitutional Symptoms (e.g., weakness, fatigue, weight loss, low hemoglobin/blood loss), Women’s Health and Gynecological Issues (e.g., menstrual problems, pregnancy, family planning), Acute Illnesses and Infections (e.g., fever, cough, cold, typhoid, dengue), and Preventive Care and Diagnostics (e.g., routine check-ups, blood tests, ultrasounds). All audio recordings were captured using the built-in microphone of a mobile device. The recording format is WAV using $16$-bit PCM at 1616kHz sampling rate.
@@ 3.2 Annotation
All recordings were processed through a structured, multi-stage annotation pipeline prior to dataset finalization. We performed a multi-stage quality check to ensure quality of the audio files, accuracy of the time-markers, as well as the correctness of the ASR transcripts. Further, the clinical summary is given by an expert clinician (Doctor) who had access to the audio and ground-truth rich transcriptions. The workflow was organized into distinct segmentation, transcription, and quality assurance phases to ensure traceability and reproducibility. We include the details of the dataset, pre-processing steps, and quality-check in the supplementary material.
@@ 3.3 Development and Evaluation set
For Tracks 1 (speaker diarization) and 2 (ASR), the data is split into $25$ hours for development and $10$ hours for blind evaluation. The development data was released in two phases: Dev Set 1 ($15$ hours) and Dev Set 2 ($10$ hours). For Tracks 3 (topic identification) and Track 4 (dialogue summarization), a separate curated subset of about $15$ hours is provided, with $10$ hours for development and $5$ hours for evaluation.
Each track has a separate set of reference annotations. The output hypothesis generated by the challenge participants were instructed to follow a specific format (like RTTM, json) for each task in order to enable the evaluation in the open-leaderboard [^3] https://www.codabench.org/competitions/13833/?secret_key=1b714e64-0f0d-4e0f-8a3c-be9b3d10f00c#.
@@ 4 Challenge Tracks and Metrics
    Track 3 - Topic Identification (TI) aims to identify the underlying topics discussed between the NPHW and the HS. This is similar to topic modeling in NLP applications. The performance metric is the Rouge-1 (R-1) and Rouge-L (R-L) [lin-2004-rouge]. Specifically, R-1 measures the overlap of individual clinical keywords (unigrams), while R-L identifies the longest common subsequence to account for sentence structure and word order.
@@ 5.1 Track-1: Speaker Diarization (SD)
The speaker diarization system is the base variant of the DiariZen [^4] https://github.com/BUTSpeechFIT/DiariZen model [han2025leveraging, han2025fine, han2025efficient], which closely follows the pyannote [bredin2023pyannote] pipeline. The pipeline follows a two-stage process, in which an end-to-end neural diarization (EEND) model first computes the local diarization results on segments of $8$s duration, followed by an Agglomerative Hierarchical Clustering (AHC) stage, where the segment-wise speaker embeddings are grouped to produce the final diarization output. DiariZen uses a custom architecture for the EEND model trained using the powerset loss [plaquet2023powerset], and ResNet34-LM trained on the VoxCeleb2 dataset [chung2018voxceleb2] to extract the speaker embeddings. AHC is carried out with the number of clusters set to $2$, the minimum cluster size set to $13$, and an AHC threshold of $0.6$. This system is refered to as Baseline-1.
For the DISPLACE-M baseline, DiariZen was evaluated under two setups—zero-shot inference and supervised fine-tuning with the development data. The fine-tuning was done with the AHC threshold set to $0.7$ and minimum cluster size set to $30$. This system is referred to as Baseline-2.
@@ 5.2 Track-2: Automatic Speech Recognition (ASR)
Similar to Track-1, both the models were evaluated with and with-out fine-tuning on the development set.
@@ 5.3 Track-3: Topic Identification (TI)
We employ an ASR–LLM cascade pipeline to identify medical topics in healthcare conversations. This approach first converts spontaneous speech into text using ASR systems, followed by an LLM, medgemma-1.5-4b-it [sellergren2025medgemma] prompting (described in the supplementary material) to generate the relevant health topics.
@@ 5.4 Track 4: Dialogue Summarization (DS)
Similar to the topic identification pipeline, we used the ASR transcript for the dialogue summarization (DS) task. For the summarization stage, we used LLAMA-3.2-3B model [mansha2025resource] as the text based summarization model in a zero-shot setting. The prompt used for summarization is described in the supplementary material.
@@ 6 Challenge Results
We also evaluated state-of-the-art closed-source models as reference systems. In particular, we evaluated Gemini 2.5 Pro [comanici2025gemini] across all four tracks and Sarvam AI’s Saaras v3777https://dashboard.sarvam.ai/speech-to-text for Track 1 (Speaker Diarization) and Track 2 (Automatic Speech Recognition).
@@ 6.1 Track 1: Speaker Diarization
Figure 2 illustrates the Eval Phase-1 results for the speaker diarization (SD) track. A total of $9$ teams participated in Phase-1 evaluation. The top four systems achieved moderate improvements over Baseline-2, as indicated by the green bars in the figure.
Figure 2: Comparison of Diarization Error Rate (DER %) for Track 1 Evaluation Phase 1 across submitted teams
@@ 6.2 Track 2: ASR
The ASR track (Track-2) received $4$ submissions, of which two teams outperformed the baseline system. Table 1 presents the Eval Phase-1 results for the ASR track. Among the baselines described in Section 5.2, fine-tuning significantly improved performance over the zero-shot setting. The fine-tuned IndicConformer achieved a tcpWER of 20.2320.23% compared to 26.7826.78% in the zero-shot setup, indicating the importance of in-domain adaptation.
@@
Table 1: ASR Track (Track-2) Results on Eval Phase-1; ZS = Zero-shot inference, FT = Fine-tuned on challenge dev-set
@@ 6.3 Track 3: Topic Identification
Table 2 presents the results for both Track $3$ and Track $4$. Track $3$ received two submissions in Eval Phase-1 and both the teams outperformed the baseline system described in Sec. 5.4.
Table 2: Results for Track-3 and Track-4.
@@ 6.4 Track 4: Dialogue Summarization
Track $4$ received three submissions in Eval Phase-1, of which two teams outperformed the baseline system described in Sec. 5.3. The top-performing team (T1) proposed an end-to-end pipeline combining VAD [Silero_VAD], speech translation and LLM-based summarization. Audio was segmented using Silero VAD, translated from Hindi to English using AudioX-North Whisper [jiviai_audiox_north_2024] and summarized with GPT-4o-mini [achiam2023gpt] using a structured clinical prompt. Team 2 (T2) used the Gemini 3 Pro end-to-end model in a 6-shot configuration, where raw audio was directly provided as input without an intermediate ASR step and the model was guided by six example transcript–summary pairs from the challenge development set.
@@ 7 Discussion and Conclusion
The DISPLACE-M challenge aims to study and investigate the difficulties that state-of-the-art systems face in transcribing and understanding front-line health conversations. Our results indicate that upstream tasks such as speaker diarization and automatic speech recognition still require work to enable seamless understanding of these conversations. The best ASR system achieves a WER of $18.2$ % (and tcpWER of $18.65$ %). Further, downstream tasks such as topic identification and dialogue summarization remain significantly more challenging due to their domain-specific nature and the complexity of interactions. Among the evaluated tasks, dialogue summarization is the most challenging and we observe that even large closed-source models struggle to generate reliable and clinically accurate medical summaries. One key reason is the conversational structure of health worker-seeker interactions, as they often contain implicit symptoms, fragmented descriptions and context-dependent dialogue, which require deeper reasoning to interpret correctly. Generating accurate summaries requires domain knowledge and multi-step reasoning to interpret symptoms and identify medically relevant information. Since the timeline was only about $6$ weeks, we have initiated a Phase-II plan for this evaluation which allows $10$ more weeks of open-leaderboard style evaluation. The Phase-II evaluation will include more languages (beyond Hindi) and will likely attract more participation.
In summary, we designed a benchmark to advance conversational AI for frontline health dialogue understanding. The evaluation framework consists of four tracks - Speaker Diarization, Automatic Speech Recognition, Topic identification and Dialogue Summarization, allowing a comprehensive evaluation pipeline from speech processing to higher-level conversational understanding. We also described the data collection and annotation process, as well as baseline systems for the proposed tasks. Further, the paper summarized the key findings from the Phase-I evaluation.

### 04 Mar 2026 v2

@@ Abstract
The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge introduces a conversational AI benchmark focused on understanding goal-oriented, real-world medical dialogues collected in the field. The challenge addresses multi-speaker interactions between healthcare workers and seekers, characterized by spontaneous, noisy and overlapping speech across Indian languages and dialects. As part of the challenge, medical conversational dataset comprising 25 hours of development data and 10 hours of blind evaluation recordings was released. We provided baseline systems within a unified end-to-end pipeline across 4 tasks - speaker diarization, automatic speech recognition, topic identification and dialogue summarization - to enable consistent benchmarking. System performance is evaluated using established metrics such as diarization error rate (DER), time-constrained minimum-permutation word error rate (tcpWER), and ROUGE-L. During this evaluation (Phase-I), $12$ teams, across the globe, actively participated pushing the baseline systems on these metrics. However, even with a 686-8 week dedicated effort from various participants, the task is shown to be substantially challenging, and the existing systems are significantly short of healthcare deployment readiness.
@@ 1 Introduction
Speech processing tools for the understanding of conversations from the healthcare domain can have transformative effects on public health systems, especially in community healthcare settings. Existing speech datasets in this domain are mostly recorded in controlled environments such as hospitals, involve structured clinician-patient conversations, and are predominantly in English. AI-driven tools developed using such datasets often fail to perform when deployed in real-world setups involving community healthcare workers. Further, there is a lack of research in low-resource languages, in particular in Indian languages. To catalyze research efforts in this space, we collected and released an annotated corpus of conversations between community health workers and healthcare seekers in informal settings in Hindi. Using the data, we hosted an open leaderboard-style challenge - ``The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge''.
@@ 2 Related Work
In the Indian context, there has been a growing interest in developing ASR systems for low-resource and multilingual settings. However, existing datasets are restricted to isolated utterances or short conversational clips [Bhogale2022EffectivenessOM, bhogale23_interspeech, Nithya2023SPRINGINXAM, javed-etal-2024-indicvoices], which limits their suitability for studying long-form, dialogue centric ASR systems. Further, there are no public datasets covering conversations in the healthcare domain.
@@
The Phase-I of the DISPLACE-M evaluations was organized between January 1, 2026 and February 27, 2026. Timeline is shown in Figure 1 . The details of the dataset (Section 3), tracks and metrics (Section 4), baseline systems (Section 5) and challenge results (Section 6) are described below.
@@ 3.1 Data Collection
We recorded around $50$ hours of annotated healthcare conversational speech from rural and semi-urban areas, collected through field deployments in the Indian states of Haryana and Bihar. Data collection was conducted by $80$ frontline community health workers across $20$ administrative blocks spanning $10$ districts. The primary participants were NPHWs (non-physician health workers) - ASHA [^1] ASHA (Accredited Social Health Activists) workers, who are government-appointed female community health workers serving as the first point of contact for healthcare seekers (HS) and Anganwadi Sevikas  [^2] Anganwadi sevikas are frontline workers under India’s Integrated Child Development Services (ICDS) program. Recordings captured natural NPHW-participant conversations occurring during routine community healthcare interactions. These interactions took place in real care environments such as Primary Health Centres (PHCs), Anganwadi centers, primary schools, open village meeting spaces (chaupals), household visits and community outreach settings. In our recordings, Hindi served as the primary language with code-switching to Indian English and with heavy use of regional dialects, including Haryanvi, Bhojpuri, and Magahi, reflecting natural linguistic practices.
The recorded conversations include speakers from different age groups. Each recording ranges from 55-$30$ minutes, typically involving goal-oriented conversations between two participants - NPHW and a healthcare seeker. The conversations, which include nature field ambiance, primarily focused on topics such as General Health and Constitutional Symptoms (e.g., weakness, fatigue, weight loss, low hemoglobin/blood loss), Women’s Health and Gynecological Issues (e.g., menstrual problems, pregnancy, family planning), Acute Illnesses and Infections (e.g., fever, cough, cold, typhoid, dengue), and Preventive Care and Diagnostics (e.g., routine check-ups, blood tests, ultrasounds). All audio recordings were captured using the built-in microphone of a mobile device. The recording format was WAV using $16$-bit PCM at 1616kHz sampling rate.
@@ 3.2 Annotation
All recordings were processed through a structured, multi-stage annotation pipeline prior to dataset finalization. We performed a multi-stage check to ensure quality of the audio files, accuracy of the time-markers and the correctness of the ASR transcripts. Further, the clinical summary is given by an expert clinician (Doctor) who had access to the audio and ground-truth rich transcriptions. The workflow was organized into distinct segmentation, transcription, and quality assurance phases to ensure traceability and reproducibility. We include the details of the dataset, pre-processing steps, and quality-checks in the supplementary material.
@@ 3.3 Development and Evaluation set
For Tracks 1 (speaker diarization) and 2 (ASR), the dataset is split into $25$ hours for development and $10$ hours for blind evaluation. The development data was released in two phases: Dev Set 1 ($15$ hours) and Dev Set 2 ($10$ hours). For Tracks 3 (topic identification) and Track 4 (dialogue summarization), a separate curated subset of about $15$ hours is provided, with $10$ hours for development and $5$ hours for evaluation.
Each track has a separate set of reference annotations. The output hypotheses generated by the challenge participants were instructed to follow a specific format (like RTTM, json) for each task in order to enable the evaluation in the open-leaderboard [^3] https://www.codabench.org/competitions/13833/?secret_key=1b714e64-0f0d-4e0f-8a3c-be9b3d10f00c#.
@@ 4 Challenge Tracks and Metrics
    Track 3 - Topic Identification (TI) aims to identify the underlying topics discussed between the NPHW and the HS. This is similar to topic modeling in NLP applications. The performance metrics are Rouge-1 (R-1) and Rouge-L (R-L) [lin-2004-rouge]. Specifically, R-1 measures the overlap of individual clinical keywords (unigrams), while R-L identifies the longest common subsequence to account for sentence structure and word order.
@@ 5.1 Track-1: Speaker Diarization (SD)
The speaker diarization system is the base variant of the DiariZen [^4] https://github.com/BUTSpeechFIT/DiariZen model [han2025leveraging, han2025fine, han2025efficient], which closely follows the pyannote [bredin2023pyannote] pipeline. The pipeline follows a two-stage process, in which an end-to-end neural diarization (EEND) model first computes the local diarization results on segments of $8$s duration, followed by an Agglomerative Hierarchical Clustering (AHC) stage, where the segment-wise speaker embeddings are grouped to produce the final diarization output. DiariZen uses a custom architecture for the EEND model trained using the powerset loss [plaquet2023powerset], and ResNet34-LM trained on the VoxCeleb2 dataset [chung2018voxceleb2] to extract the speaker embeddings. AHC is carried out with the number of clusters set to $2$, the minimum cluster size set to $13$, and an AHC threshold of $0.6$. This system is referred to as Baseline-1.
For the DISPLACE-M baseline, DiariZen was evaluated under two setups - zero-shot inference and supervised fine-tuning with the development data. The fine-tuning was done with the AHC threshold set to $0.7$ and minimum cluster size set to $30$. This system is referred to as Baseline-2.
@@ 5.2 Track-2: Automatic Speech Recognition (ASR)
Similar to Track-1, both the models were evaluated with and without fine-tuning on the development set.
@@ 5.3 Track-3: Topic Identification (TI)
We employ an ASR–LLM cascade pipeline to identify medical topics in healthcare conversations. This approach first converts spontaneous speech into text using ASR systems, followed by an LLM, medgemma-1.5-4b-it [sellergren2025medgemma] prompting to generate the relevant health topics.
@@ 5.4 Track 4: Dialogue Summarization (DS)
Similar to the topic identification pipeline, we used the ASR transcript for the dialogue summarization (DS) task. For the summarization stage, we used LLAMA-3.2-3B model [mansha2025resource] as the text-based summarization model in a zero-shot setting. The prompt used for summarization is described in the supplementary material.
@@ 6 Challenge Results
We also benchmarked state-of-the-art closed-source models as reference systems. In particular, we evaluated Gemini 2.5 Pro [comanici2025gemini] across all four tracks and Sarvam AI’s Saaras v3777https://dashboard.sarvam.ai/speech-to-text for Track 1 (Speaker Diarization) and Track 2 (Automatic Speech Recognition).
@@ 6.1 Track 1: Speaker Diarization
Figure 3 illustrates the Eval Phase-I results for the SD track. A total of $9$ teams participated in Phase-I evaluation. The top four systems achieved moderate improvements over Baseline-2, as indicated by the green bars in the figure.
Figure 3: Comparison of Diarization Error Rate (DER %) for Track 1 Evaluation Phase-I across submitted teams.
@@ 6.2 Track 2: ASR
The ASR track (Track-2) received $4$ submissions, of which two teams outperformed the baseline system. Table 1 presents the Eval Phase-I results for the ASR track. Among the baselines described in Section 5.2, fine-tuning significantly improved performance over the zero-shot setting. The fine-tuned IndicConformer achieved a tcpWER of 20.2320.23% compared to 26.7826.78% in the zero-shot setup, indicating the importance of in-domain adaptation.
@@
Table 1: ASR Track (Track-2) Results on Eval Phase-I; ZS = Zero-shot inference, FT = Fine-tuned on challenge dev-set.
@@ 6.3 Track 3: Topic Identification
Table 2 presents the results for both Track $3$ and Track $4$. Track $3$ received two submissions in Eval Phase-I and both teams outperformed the baseline system described in Sec. 5.4.
Table 2: Results for Track-3 (TI) and Track-4 (DS).
@@ 6.4 Track 4: Dialogue Summarization
Track $4$ received three submissions in Eval Phase-I, of which two teams outperformed the baseline system. The top-performing team (T1) proposed an end-to-end pipeline combining Silero-VAD [Silero_VAD] for segmentation, speech translation from Hindi to English using AudioX-North Whisper [jiviai_audiox_north_2024] and GPT-4o-mini-based summarization using a structured clinical prompt. Team 2 (T2) used the Gemini-3 Pro end-to-end model in a $6$-shot configuration, where raw audio was directly provided as input without an intermediate ASR step and the model was guided by six example transcript–summary pairs from the challenge development set.
@@ 7 Discussion and Conclusion
The DISPLACE-M challenge aims to study and investigate the difficulties that state-of-the-art systems face in transcribing and understanding frontline health conversations. Our results indicate that upstream tasks such as speaker diarization and automatic speech recognition still requires work to enable seamless understanding of these conversations. The best ASR system achieves a WER of $18.2$ % (and tcpWER of $18.65$ %). Further, downstream tasks such as topic identification and dialogue summarization remain significantly more challenging due to their domain-specific nature and the complexity of interactions. Among the evaluated tasks, dialogue summarization is the most challenging and we observe that even large closed-source models struggle to generate reliable and clinically accurate medical summaries. One key reason is the conversational structure of health worker-seeker interactions, as they often contain implicit symptoms, fragmented descriptions and context-dependent dialogue, which require deeper reasoning to interpret correctly. Generating accurate summaries requires domain knowledge and multi-step reasoning to interpret symptoms and identify medically relevant information. Since the timeline was only about $6$ weeks, we have initiated a Phase-II plan for this evaluation which allows $10$ more weeks of open-leaderboard style evaluation. The Phase-II evaluation will include more languages (beyond Hindi) and will likely attract more participation.
In summary, we designed a benchmark to advance conversational AI for frontline health dialogue understanding. The evaluation framework consists of four tracks - Speaker Diarization, Automatic Speech Recognition, Topic identification and Dialogue Summarization, allowing a comprehensive evaluation pipeline from speech processing to higher-level conversational understanding. We also described the data collection and annotation process, as well as baseline systems for the proposed tasks. Further, the paper summarizes the key findings from the Phase-I evaluation.